A 0.100 M solution of an unknown acid, HA, is 1.47
Answer #1

HA dissociates as: - - - - - > H+ + A- o Initial Change Equilibrium HA 0.1 -X 0.1-x of x * *

1)

% dissociation = (x*100)/c

1.47= x*100/0.1

x = 0.0015

HA dissociates as:

HA -----> H+ + A-

Initial 0.1 0 0

Change -x +x +x

Equilibrium 0.1-x x x

Ka = [H+][A-]/[HA]

Ka = x*x/(c-x)

Ka = 1.47*10^-3*1.47*10^-3/(0.1-1.47*10^-3)

Ka = 2.193*10^-5

Answer: 2.19E-5

2)

[H+] = x = 1.47*10^-3 M

use:

pH = -log [H+]

= -log (1.47*10^-3)

= 2.8327

Answer: 2.83